Problem 4. (8 points) We deal from a well-shuffled 52-card deck. What is the probability that the 13th card is the first king to be dealt? Use the Multiplication Rule to solve the problem.
Answer #1

P(13 th card is the first king)

=P(there is no king in first 12 card)*P(13th card is king|no king in first 12 cards)

=P(select 12 card from 48 non king cards)*P(selecting a king from 40 cards after removal of 12 cards)

=4852 1212(4/40) =0.33757503*0.1=0.03376